OpenStudy (anonymous):

can someone help me understand this question: Which statements are true? Choose all answers that are correct. A. 7 + (–10) < –13 + 17 B. 22 + (–13) > 18 + (–7) C. –18 + 24 = –23 + 17 D. –11 + (–8) > 4 + (–26)

OpenStudy (igreen):

It wants you to solve both sides of the sign, and see if they are true.

OpenStudy (igreen):

7 + (–10) < –13 + 17 Simplify: 7 - 10 < -13 + 17 Subtract: -3 < -13 + 17 Add: -3 < 4 This is a true statement because -3 is LESS THAN 4.

OpenStudy (igreen):

22 + (–13) > 18 + (–7) Simplify: 22 - 13 > 18 - 7 Subtract: 9 > 11 This is a false statement because 9 is not greater than 11.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

–18 + 24 = –23 + 17 Add: 6 = -23 + 17 Add: 6 = -6 This is a false statement because 6 does not equal -6.

OpenStudy (igreen):

–11 + (–8) > 4 + (–26) Simplify: -11 - 8 > 4 - 26 Subtract: -19 > 4 - 26 Subtract: -19 > -22 This is a true statement because 19 is greater than -22.

OpenStudy (igreen):

So your answers are A and D.

OpenStudy (anonymous):

kk thx i was actaully figuring it out as i was going

OpenStudy (igreen):

Okay, no problem.